what is the concentration in mol dm of a solution of potassium hydroxide, KOH containing 4,13 gdm
1
Expert's answer
2012-11-12T10:58:55-0500
n = 4,13 / 56 = 0.07375 mol
Cm = n/V
Cm = 0.07375 / 1 = 0.07375 mol/dm3